# Scale-Aware Pre-Training for Human-Centric Visual Perception: Enabling Lightweight and Generalizable Models

### v1

@@ Abstract
corresponding author.
Human-centric visual perception (HVP) has recently achieved remarkable progress due to advancements in large-scale self-supervised pretraining (SSP). However, existing HVP models face limitations in adapting to real-world applications, which require general visual patterns for downstream tasks while maintaining computationally sustainable costs to ensure compatibility with edge devices. These limitations primarily arise from two issues: 1) the pretraining objectives focus solely on specific visual patterns, limiting the generalizability of the learned patterns for diverse downstream tasks; and 2) HVP models often exhibit excessively large model sizes, making them incompatible with real-world applications. To address these limitations, we introduce Scale-Aware Image Pretraining (SAIP), a novel SSP framework enabling lightweight vision models to acquire general patterns for HVP. Specifically, SAIP incorporates three learning objectives based on the principle of cross-scale consistency: 1) Cross-scale Matching (CSM) which contrastively learns image-level invariant patterns from multi-scale single-person images; 2) Cross-scale Reconstruction (CSR) which learns pixel-level consistent visual structures from multi-scale masked single-person images; and 3) Cross-scale Search (CSS) which learns to capture diverse patterns from multi-scale multi-person images. Three objectives complement one another, enabling lightweight models to learn multi-scale generalizable patterns essential for HVP downstream tasks. Extensive experiments conducted across 12 HVP datasets demonstrate that SAIP exhibits remarkable generalization capabilities across 9 human-centric vision tasks. Moreover, it achieves significant performance improvements over existing methods, with gains of 3%-13% in single-person discrimination tasks, 1%–11% in dense prediction tasks, and 1%–6% in multi-person visual understanding tasks. All code and checkpoints will be publicly available [^1] https://github.com/stoa-xh91/SAIPv1.
@@
1 Introduction
With the advances in large-scale self-supervised pretraining (SPP) [1, 14, 9, 36, 32, 33], remarkable progress has been made in human-centric visual perception (HVP). Numerous general HVP models [42, 16, 35, 2, 4] have been introduced, substantially advancing the state of the art.
Despite the differences in architectures, existing general HVP models follow a common training paradigm: they utilize self-supervised pretraining to enable vision models to learn generalizable visual patterns for downstream tasks, followed by fine-tuning the pretrained models to adapt to specific task demands. Based on this, existing works aim to study: 1) how to collect high-quality person images that benefit model training [35, 23]; and 2) how to design a proper pretext task that facilitates the learning of human visual patterns [42, 16, 2]. However, all of these works ignore limitations in adapting to real-world applications, which require general visual patterns for downstream tasks while maintaining computationally sustainable costs to ensure compatibility with edge devices. Such limitations primarily arise from two drawbacks:
(1) Specific visual patterns. As shown in Fig. 1(a), human-centric visual perception in real-world scenes often involves numerous instance discrimination tasks and diverse dense prediction tasks, which require general human visual patterns capable of perceiving multiple persons across a broad range of scales. However, existing methods [42, 2, 16, 32, 1, 14] pretrain vision models by learning specific visual patterns, i.e. single instance at a single scale. When directly applying these methods to lightweight vision models, a significant misalignment arises between the learned visual patterns and those required for generalization to downstream tasks. As illustrated in Fig. 2, when faced with diverse scales of one instance, vision models pretrained using existing methods fail to capture consistent human visual patterns. Additionally, when faced with multiple instances in one image, captured visual patterns appear to focus on the central instance while neglecting the others. This phenomenon suggests that these pretrained models are prone to overfitting specific human visual patterns, i.e. single instance at single scale, thus limiting their generalizability to a wide range of downstream tasks.
@@ 1 Introduction
Aforementioned limitations motivate us to rethink a significant question that is considerably less studied: is it possible to design a pretraining method that enables a HVP model to be generalizable for diverse downstream tasks while maintaining lightweight, and how can this be achieved ? In this work, we seek to answer this question by proposing Scale-Aware Image Pretraining (SAIP), a novel self-supervised pretraining framework particularly designed for lightweight HVP models. Specifically, deep neural architectures with around 5 millions parameters are utilized as the lightweight vision models. Within the SAIP framework, those lightweight vision models are guided to learn cross-scale consistent patterns under both single-instance and multi-instance settings, aiming to align the learned visual patterns with the requirement of cross-scale consistency in real-world scenes. To this end, three learning tasks are proposed: 1) Cross-scale Matching (CSM) which contrastively learns image-level invariant patterns from multi-scale person-centric images; 2) Cross-scale Reconstruction (CSR) which learns pixel-level consistent patterns from cross-scale masked images; and 3) Cross-scale Search (CSS) which learns to capture diverse patterns from multi-person images across multiple scales. Three tasks complement one another, enabling lightweight models to effectively capture general human vision patterns required by diverse HVP downstream tasks. To summary, the contributions of this paper are three fold.
@@
  • Flexibility and strong generalization. SAIP is model-agnostic, allowing it to be applied for any model architecture. Based on SAIP, we have successfully established generalizable HVP models with three lightweight architectures (i.e. ViT, ConvNet and the hybrid), and resulting models are comprehensively evaluated across 12 human-centric visual perception datasets. Specifically, SAIP surpasses previous SSP methods across 9 specific human-centric tasks by a large margin. It achieves significant improvements in single-person discrimination tasks by 3%-13%, dense predictions by 1%-11%, and multi-person visual understanding by 1%-6%, exhibiting a remarkable generalization capability.
@@
2 Related Work
Human-centric Visual Perception. The intelligent understanding of humans through visual media, such as image and video, has been a long-standing topic in the computer vision community. Optimally addressing human-centric visual perception would significantly support a wide range of visual applications, such as person ReID in security system, human parsing in metaverse, and person detection in autonomous driving. In general, HVP tasks can be categorized into three groups: 1) Single-person discrimination (e.g. Person ReID or Pedestrian Attribute Recognition) [17, 15, 19], which focuses on the identity information of a person; 2) Single-person dense predictions task (e.g. pose estimation or body-part segmentation) [40, 41, 24, 26], which particularly focuses on the shape or posture of a person; and 3) multi-person visual understanding (e.g. multiple human parsing or part-level attribute parsing) [41, 37, 3, 12, 18], which is a hybrid task that simultaneously perform person detection and single-person recognition in crowded scene. These tasks are highly correlated as they share a common goal of visually understanding human body structure. Consequently, recent works [35, 16, 4, 21] attempt to unify these tasks by pretraining a large model on extensive publicly available datasets with multi-task labels. However, the scarcity of high-quality labels significantly limits the scalability of the pretrained models.
@@ 2 Related Work
Self-supervised Visual Pretraining. Learning meaningful patterns from massive unlabeled data has been shown to be a promising approach for developing visual foundation models. As presented in Fig 1.(a), mainstream studies in general can be categorized into two categories: contrastive learning (CL) [1, 13, 16, 2] and masked image modeling (MIM) [14, 9, 36, 33, 42]. The CL models image similarity and dissimilarity between two or more views of the same object, aiming to capture long-range global patterns such as shape. In contrast, the MIM focuses on reconstruction of masked images, preferring high-frequency signals that are texture-oriented. Recent studies [32] reveal that CL and MIM can complement each other and their straightforward combination can enhance model pretraining.
@@ 3 Method
3.1 Model Architecture
(1) Image encoder, a visual feature extractor designed to tokenize an image into numerous visual tokens. The image encoder can be built on various architectures, such as vision transformer (ViT), convolution neural network (ConvNet) or even a hybrid of ViT and ConvNet. In this work, we adopt plain vision transformer [8], StarNet [31] and TinyViT [39] as the instantiations respectively for ViT-, ConvNet- and hybrid-based image encoder. Formally, given an image $I\in\mathbb{R}^{3\times H\times W}$, the image encoder outputs an image token $F\in\mathbb{R}^{D}$ representing entire image and $L$ patch tokens $F^{\textquotesingle}\in\mathbb{R}^{L\times D}$ as well.
(3) Task decoders, which utilize image-level decoder or pixel-level decoder to perform three proxy tasks. In particular, the image-level decoder is a stack of feed-forward layers that project the image token $F$ to instance representation $\hat{F}\in\mathbb{R}^{d}$ for the purpose of cross-scale matching. The pixel-level decoder, denoted as $\phi([query],[key],[value])$, comprises a stack of interactive computation blocks, each of which sequentially includes self-attention layer, cross-attention layer and feed-forward layer. Two types of image pair is considered for the pixel-level decoder: 1) masked image and non-masked image; and 2) instance image and multi-instances image. The pixel-level decoder take the former pair as input for cross-scale reconstruction task, and the latter pair for cross-scale search.
3.2 Pretraining Objectives
We optimize lightweight models using three learning objectives. To facilitate the pretraining process, we prepare three types of inputs: an anchor image $I_{0}$ of a person, $N$ images with different resolutions for the person $\{I_{t}\}\sum_{t=1}^{N}$, and a multi-person image $I_{N+1}$. Specifically, the multi-resolution images are obtained by randomly resizing the anchor image to different scales ranging from 0.75 to 1.5, while the multi-person image is synthesized by applying a simple copy-paste technique [11] to multi-resolution images. Based on these different inputs, task decoders are activated to compute the three losses, as delineated below.
$\begin{array}[]{lll}\ell_{csm}&=&\frac{1}{N}\sum\limits_{t}^{N}-\hat{F_{0}}log(\hat{F_{t}})\\ \end{array}\vspace*{-0.1in}$ (1)
(2) Cross-scale Reconstruction (CSR) designed to learn common structure of a person instance at different scales. To achieve this, we utilize the pixel-level decoder with cross-attention to reconstruct a masked image conditioned on anchor image. Specifically, we randomly select one of multi-resolution images and divide it into non-overlapping patches. A subset of these patches (e.g., mask ratio is 75% in general) is randomly masked, leaving the rest visible. Regarding the masked image as the query, the pixel-level decoder is activated to project patch tokens of anchor image to RGB space for the reconstruction of the masked image. The similarity between the predicted image and selected image is penalized through mean square errors as formulated in Equ 2:
$\begin{array}[]{lll}\ell_{css}&=&\sum_{t=1}^{N}\left\|\tau(\phi(F_{t}^{\textquotesingle},F_{N+1}^{\textquotesingle},F_{N+1}^{\textquotesingle}))-M_{t}\right\|^{2}_{2}\\ \end{array}\vspace*{-0.1in}$ (3)
$\begin{array}[]{lll}&\mathop{min}\limits_{\theta}\mathbb{E}[\mathcal{L}(I_{0},I_{t},I_{N+1})],&\\ &s.t.,\,\,\,\mathcal{L}=\ell_{csm}+\ell_{csr}+\ell_{css}&\end{array}\vspace*{-0.1in}$ (4)
where $\theta$ is the learnable parameters of a lightweight model.
4 Experiments
In this section, we first present an overview of the experimental settings, followed by the implementation details. Subsequently, we conduct comprehensive ablation study to explore major properties of proposed SAIP. Finally we analyze the merit of SAIP through a comprehensive comparison with existing methods on 9 human-centric tasks, which include image-to-image person ReID (I2I ReID), text-to-image person ReID (T2I ReID), person attribute recognition (PAR), human pose estimation (HPE), person landmarks detection (PLD), body part segmentation (BPS), pedestrian detection (PD), multiple human parsing (MHP) and part-level attribute parsing (PAP).
4.1 Experimental Settings
4.2 Implementation Details
Pretraining. All lightweight models are pretrained or finetuned using 8 A6000 48G GPUs. We use the AdamW [30] as the optimizer and set the effective batch size is 2048 (i.e. 256 per GPU). Each model is pretrained from scratch for 300 epochs in default, with a initial learning rate of 2.5e-4, which is declined by Cosine Annealing scheduler [29]. Other settings are identical to that of DINO [1].
@@ 4.2 Implementation Details
Table 1: Quantitative comparison with state-of-the-art pretraining methods on 6 single-person discrimination datasets. * denotes the expert encoder in SAIP framework is a pretrained big model such as PATH [35]. $\uparrow$ means the larger value the better performance.
4.3 Main Results
@@
Table 2: Quantitative comparison with state-of-the-art pretraining methods on 3 single-person dense predictions datasets.

### v2

@@ Abstract
Human-centric visual perception (HVP) has recently achieved remarkable progress due to advancements in large-scale self-supervised pretraining (SSP). However, existing HVP models face limitations in adapting to real-world applications, which require general visual patterns for downstream tasks while maintaining computationally sustainable costs to ensure compatibility with edge devices. These limitations primarily arise from two issues: 1) the pretraining objectives focus solely on specific visual patterns, limiting the generalizability of the learned patterns for diverse downstream tasks; and 2) HVP models often exhibit excessively large model sizes, making them incompatible with real-world applications. To address these limitations, we introduce Scale-Aware Image Pretraining (SAIP), a novel SSP framework pretraining lightweight vision models to acquire general patterns for HVP. Specifically, SAIP incorporates three learning objectives based on the principle of cross-scale consistency: 1) Cross-scale Matching (CSM) which contrastively learns image-level invariant patterns from multi-scale single-person images; 2) Cross-scale Reconstruction (CSR) which learns pixel-level consistent visual structures from multi-scale masked single-person images; and 3) Cross-scale Search (CSS) which learns to capture diverse patterns from multi-scale multi-person images. Three objectives complement one another, enabling lightweight models to learn multi-scale generalizable patterns essential for HVP downstream tasks. Extensive experiments conducted across 12 HVP datasets demonstrate that SAIP exhibits remarkable generalization capabilities across 9 human-centric vision tasks. Moreover, it achieves significant performance improvements over existing methods, with gains of 3%-13% in single-person discrimination tasks, 1%–11% in dense prediction tasks, and 1%–6% in multi-person visual understanding tasks. Code and models are released [^1] https://github.com/stoa-xh91/SAIPv1 for research purpose.
Index Terms:
Article submission, IEEE, IEEEtran, journal, , paper, template, typesetting.
I Introduction
@@
With the advances in large-scale self-supervised pretraining (SPP) [1, 2, 3, 4, 5, 6, 7], remarkable progress has been made in human-centric visual perception (HVP). Numerous general HVP models [8, 9, 10, 11, 12] have been introduced, substantially advancing the state of the art. Despite the differences in architectures, existing general HVP models follow a common training paradigm: they utilize self-supervised pretraining to enable vision models to learn generalizable visual patterns for downstream tasks, followed by fine-tuning the pretrained models to adapt to specific task demands. Based on this, existing works aim to study: 1) how to collect high-quality person images that benefit model training [10, 13]; and 2) how to design a proper pretext task that facilitates the learning of human visual patterns [8, 9, 11]. However, all of these works ignore limitations in adapting to real-world applications, which require general visual patterns for downstream tasks, while maintaining a lightweight model architecture with computationally sustainable costs to ensure edge-device compatibility. Such limitations primarily arise from two drawbacks:
(1) Specific visual patterns. In general, human-centric visual perception in real-world scenes often involves numerous instance discrimination tasks and diverse dense prediction tasks, which require general human visual patterns capable of perceiving multiple persons across a broad range of scales. However, existing methods [8, 11, 9, 5, 1, 2] pretrain vision models by learning specific visual patterns, i.e., single instance at a single scale. When directly applying these methods to lightweight vision models, a significant misalignment arises between the learned visual patterns and those required for generalization to downstream tasks. As illustrated in Fig. 1, when faced with diverse scales of one instance, vision models pretrained using existing methods fail to capture consistent human visual patterns. Additionally, when faced with multiple instances in one image, captured visual patterns appear to focus on the central instance while neglecting the others. This phenomenon suggests that these pretrained models are prone to overfitting specific human visual patterns, i.e., single instance at single scale, thus limiting their generalizability to a wide range of downstream tasks (Fig. 2(a)).
@@
To address these limitations, a viable direction is the large-scale pretraining of lightweight model. This motivates us to rethink a significant question that is considerably less studied: is it possible to design a pretraining method that enables a lightweight HVP model to be generalizable for diverse downstream tasks, and how can this be achieved ? In this work, we seek to answer this question by proposing Scale-Aware Image Pretraining (SAIP), a novel self-supervised pretraining framework particularly designed for lightweight HVP models. As illustrated in Fig. 2(b), deep neural architectures with around 5 millions parameters are utilized as the lightweight vision models. Within the SAIP framework, those lightweight vision models are guided to learn cross-scale consistent patterns under both single-instance and multi-instance settings, aiming to align the learned visual patterns with the requirement of cross-scale consistency in real-world scenes. To this end, three learning tasks are proposed: 1) Cross-scale Matching (CSM) which contrastively learns image-level invariant patterns from multi-scale person-centric images; 2) Cross-scale Reconstruction (CSR) which learns pixel-level consistent patterns from cross-scale masked images; and 3) Cross-scale Search (CSS) which learns to capture diverse patterns from multi-person images across multiple scales. Three tasks complement one another, enabling lightweight models to effectively capture diverse vision patterns such as global discrimination pattern, local structure pattern, and regional discrimination pattern. Consequently, SAIP pretraining facilitates efficient transfer learning for various HVP downstream tasks. To summary, the contributions of this paper are three fold.
@@
  • Flexibility and strong generalization. SAIP is model-agnostic, allowing it to be applied for any model architecture. Based on SAIP, we have successfully established generalizable HVP models with three lightweight architectures (i.e., ViT, ConvNet and the hybrid), and resulting models are comprehensively evaluated across 12 human-centric visual perception datasets. Specifically, SAIP surpasses previous SSP methods across 9 specific human-centric tasks by a large margin. It achieves significant improvements in single-person discrimination tasks by 3%-13%, dense predictions by 1%-11%, and multi-person visual understanding by 1%-6%, exhibiting a remarkable generalization capability.
@@
II Related Work
Human-centric Visual Perception. The intelligent understanding of humans through visual media, such as image and video, has been a long-standing topic in the computer vision community. Optimally addressing human-centric visual perception would significantly support a wide range of visual applications, such as person ReID in security system, human parsing in metaverse, and person detection in autonomous driving. In general, HVP tasks can be categorized into three groups: 1) Single-person discrimination task (e.g., Person ReID or Pedestrian Attribute Recognition) [17, 18, 19], which focuses on the identity information of a person; 2) Single-person dense predictions task (e.g., pose estimation or body-part segmentation) [20, 21, 22, 23], which particularly focuses on the shape or posture of a person; and 3) multi-person visual understanding task (e.g., multiple human parsing or part-level attribute parsing) [21, 24, 25, 26, 27], which is a hybrid task that simultaneously perform person detection and single-person recognition in crowded scene. These tasks are highly correlated as they share a common goal of visually understanding human body structure. Consequently, recent works [10, 9, 12, 15] attempt to unify these tasks by pretraining a large model on extensive publicly available datasets with multi-task labels. However, the scarcity of high-quality labels significantly limits the scalability of the pretrained models.
@@ II Related Work
Self-supervised Visual Pretraining. Learning meaningful patterns from massive unlabeled data has been shown to be a promising approach for developing visual foundation models. As presented in Fig 2.(a), mainstream studies in general can be categorized into two paradigms: contrastive learning (CL) [1, 28, 9, 11, 29] and masked image modeling (MIM) [2, 3, 4, 7, 8]. The CL models image similarity and dissimilarity between two or more views of the same object, aiming to capture long-range global patterns such as shape. In contrast, the MIM focuses on reconstruction of masked images, preferring high-frequency signals that are texture-oriented. Recent studies [5] reveal that CL and MIM can complement each other and their straightforward combination can enhance model pretraining.
@@ III Method
Supervised Pretraining StarNet_S3 [30] 5.8M IN1k 90.5 65.8 55.4 44.8 79.1 71.9
CSPNeXt_S [31] 4.4M AIC+COCO 89.8 63.6 47.5 33.6 78.9 67.3
ViT_Tiny [32] 5.5M IN1K 89.3 65.9 51.1 45.9 74.5 68.0
EdgeNeXt [33] 5.6M IN1K 89.5 67.5 58.0 47.4 80.6 70.8
TinyViT_5M [34] 5.4M IN1K 89.1 68.9 57.8 40.9 76.5 69.5
Self-supervised Pretraining
@@
TABLE I: Quantitative comparison with state-of-the-art pretraining methods on 6 single-person discrimination datasets. * denotes the expert encoder is a pretrained model such as PATH [10]. $\uparrow$ means the larger value the better performance.
III-A Model Architecture
(1) Image encoder, a visual feature extractor designed to tokenize an image into numerous visual tokens. The image encoder can be built on various architectures, such as vision transformer (ViT), convolution neural network (CNN) or even a hybrid of ViT and CNN. In this work, we adopt plain vision transformer [32], StarNet [30] and TinyViT [34] as the instantiations respectively for ViT-, CNN- and hybrid-based image encoder. Formally, given an image $I\in\mathbb{R}^{3\times H\times W}$, the image encoder outputs an image token $F\in\mathbb{R}^{D}$ representing entire image and $L$ patch tokens $F^{\textquotesingle}\in\mathbb{R}^{L\times D}$ as well.
(3) Task decoders, which utilize image-level decoder or pixel-level decoder to perform three proxy tasks. In particular, the image-level decoder is a stack of feed-forward layers that project the image token $F$ to instance representation $\hat{F}\in\mathbb{R}^{d}$ for the purpose of cross-scale matching. The pixel-level decoder, denoted as $\phi([query],[key],[value])$, comprises a stack of interactive computation blocks, each of which sequentially includes self-attention layer, cross-attention layer and feed-forward layer. Two types of image pair is considered for the pixel-level decoder: 1) masked image and non-masked image; and 2) single-instance image and multi-instances image. The pixel-level decoder take the former pair as input for cross-scale reconstruction task, and the latter pair for cross-scale search.
III-B Pretraining Objectives
We optimize lightweight models using three learning objectives. To facilitate the pretraining process, we prepare three types of inputs: an anchor image $I_{0}$ of a person, $N$ images with different resolutions for the person $\{I_{t}\}_{t=1}^{N}$, and a multi-person image $I_{N+1}$. Specifically, the multi-resolution images are obtained by randomly resizing the anchor image to different scales ranging from 0.75 to 1.5, while the multi-person image is synthesized by applying a simple copy-paste technique [35] to multi-resolution images. Based on these different inputs, task decoders are activated to compute the three losses, as delineated below.
$\begin{array}[]{lll}\ell_{csm}&=&\frac{1}{N}\sum\limits_{t}^{N}-\hat{F_{0}}log(\hat{F_{t}})\\ \end{array}$ (1)
(2) Cross-scale Reconstruction (CSR) designed to learn common structure of a person instance at different scales. To achieve this, we utilize the pixel-level decoder with cross-attention to reconstruct a masked image conditioned on anchor image. Specifically, we randomly select one of multi-resolution images and divide it into non-overlapping patches. A subset of these patches (e.g., mask ratio is 75% in general) is randomly masked, leaving the rest visible. Regarding the masked image as the query, the pixel-level decoder is activated to project patch tokens of anchor image to RGB space for the reconstruction of the masked image. The similarity between the predicted image and selected image is measured by mean square errors as formulated in Equ 2:
$\begin{array}[]{lll}\ell_{css}&=&\sum_{t=1}^{N}\left\|\tau(\phi(F_{t}^{\textquotesingle},F_{N+1}^{\textquotesingle},F_{N+1}^{\textquotesingle}))-M_{t}\right\|^{2}_{2}\\ \end{array}$ (3)
$\begin{array}[]{lll}&\mathop{min}\limits_{\theta}\mathbb{E}[\mathcal{L}(I_{0},I_{t},I_{N+1})],&\\ &s.t.,\,\,\,\mathcal{L}=\ell_{csm}+\ell_{csr}+\ell_{css}&\end{array}$ (4)
where $\theta$ is the learnable parameters of a lightweight model.
@@ III-B Pretraining Objectives
Supervised Pretraining StarNet_S3 [30] 5.8M IN1k 71.9 75.1 50.5 62.4 52.4 61.9
CSPNeXt_S [31] 4.4M AIC+COCO 70.0 73.3 50.3 62.3 54.6 63.6
ViT_Tiny [32] 5.5M IN1K 69.3 72.6 45.0 58.1 48.9 63.5
EdgeNeXt [33] 5.6M IN1K 71.6 74.8 50.5 62.2 54.9 65.5
TinyViT_5M [34] 5.4M IN1K 69.2 72.7 53.7 65.0 56.2 67.7
Self-supervised Pretraining
@@
TABLE II: Quantitative comparison with state-of-the-art pretraining methods on 3 single-person dense predictions datasets.
IV Experiments
In this section, we first present an overview of the experimental settings, followed by the implementation details. Subsequently, we analyze the merit of SAIP through a comprehensive comparison with existing methods on 9 human-centric tasks, which include image-to-image person ReID (I2I ReID), text-to-image person ReID (T2I ReID), person attribute recognition (PAR), human pose estimation (HPE), person landmarks detection (PLD), body part segmentation (BPS), pedestrian detection (PD), multi-person human parsing (MHP), and part-level attribute parsing (PAP). Finally, we conduct comprehensive ablation study to explore major properties of proposed SAIP.
IV-A Experimental Settings
IV-B Implementation Details
Pretraining. All lightweight models are pretrained or finetuned using 8 A6000 48G GPUs. We use the AdamW [45] as the optimizer and set the effective batch size is 2048 (i.e., 256 per GPU). Each model is pretrained from scratch for 300 epochs in default, with a initial learning rate of 2.5e-4, which is declined by Cosine Annealing scheduler [46]. Other settings are identical to that of DINO [1].